A drug is being delivered into a patient’s arm at a rate of 319 mL min–1. The drug is being delivered from a syringe through a needle with an internal diameter of 1 mm. This drug has a density of 1050 kg m–3. Is the flow laminar or turbulent? (ηdrug = 8.90×10–4 Pa s)
1
Expert's answer
2021-07-25T09:16:57-0400

For the flow in a pipe (needle) the Reynold's number is what determines the laminar/turbulent flow. The Reynold's number is given as follows (see https://en.wikipedia.org/wiki/Reynolds_number#Flow_in_a_pipe):


"Re = \dfrac{\rho Q D_H}{\mu A}"

where "\rho = 1050kg/m^3" is the density of the drug, "Q = 319\space mL/min = \dfrac{319}{60}\times 10^{-6}\space m^3/s" is the volumetric flow rate, "D_H = 1mm = 10^{-3}m" is the internal diameter of the needle, "A = \pi D_H^2/4" is the cross-sectional area of the needle, and "\mu = 8.90\times 10^{-4}Pa\cdot s" is the dynamic viscosity of the drug.

Thus, obtain:


"Re = \dfrac{\rho Q D_H}{\mu\cdot \dfrac{\pi D_H^2}{4}} = \dfrac{4\rho Q}{\pi\mu D_H}\\
Re = \dfrac{4\cdot 1050\cdot \dfrac{319}{60}\times 10^{-6}}{\pi\cdot 8.90\times 10^{-4}\cdot 10^{-3}} \approx 7990"

Since "Re>2900", the flow is turbulent.


Answer. Turbulent.